how do yu solve y=4x-8 and y=2x+10
Found 2 solutions by checkley77, rapaljer:
Answer by checkley77(12844)

y=4x-8
y=2x+10 SUBTRACT.
------------------
0=2X-18
2X=18
X=18/2
X=9
Y=4*9-8
Y=36-8
Y=28 ANS.
PROOF:
28=2*9+10
28=18+10
28=28

Answer by rapaljer(4671)

Since y=y, you can set each equation equal to the other:
4x-8=2x+10

Subtract 2x from each side:
2x-8=10

Add +8 to each side:
2x=18
x=9

Since y=4x-8, substitute x=9
y=4*9-8
y=36-8=28

The solution is (9,28)

Check: Substitute both numbers into the second equation:
y = 2x+10
28=2*9+10
28=18+10 It checks!!

For additional explanation on solving Systems of Equations, please see my own website. Click on my tutor name "rapaljer" and then click on the link that is given there. From my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Select "Basic Algebra," and look for "Chapter 4". See "Section 4.07 Systems of Equations." These sections include my own explanation written especially for students who may have trouble understanding these skills, examples, exercises, and ALL the answers. Many of the exercises are solved in detail on the "Math in Living Color" pages that go with these pages.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus